# MOREL: Enhancing Adversarial Robustness through Multi-Objective Representation Learning

### v2

@@ Abstract
Extensive research has shown that deep neural networks (DNNs) are vulnerable to slight adversarial perturbations—small changes to the input data that appear insignificant but cause the model to produce drastically different outputs. In addition to augmenting training data with adversarial examples generated from a specific attack method, most of the current defense strategies necessitate modifying the original model architecture components to improve robustness or performing test-time data purification to handle adversarial attacks. In this work, we demonstrate that strong feature representation learning during training can significantly enhance the original model’s robustness. We propose MOREL, a multi-objective feature representation learning approach, encouraging classification models to produce similar features for inputs within the same class, despite perturbations. Our training method involves an embedding space where cosine similarity loss and multi-positive contrastive loss are used to align natural and adversarial features from the model encoder and ensure tight clustering. Concurrently, the classifier is motivated to achieve accurate predictions. Through extensive experiments, we demonstrate that our approach significantly enhances the robustness of DNNs against white-box and black-box adversarial attacks, outperforming other methods that similarly require no architectural changes or test-time data purification. Our code is available at https://github.com/salomonhotegni/MOREL.
@@ 1 Introduction
The deployment of deep neural networks (DNNs) in critical vision applications such as autonomous driving and medical diagnosis (Bojarski et al., 2016; Miotto et al., 2018) underscores the need for robust models capable of reliably handling real-world scenarios. However, extensive research has demonstrated that DNNs are vulnerable to adversarial examples—inputs crafted by adding imperceptible perturbations that can cause the model to make incorrect predictions with high confidence (Nguyen et al., 2015; Szegedy et al., 2013; Goodfellow et al., 2014). This vulnerability poses significant challenges to the security and reliability of AI systems, especially in safety-critical environments. To mitigate the risks posed by adversarial attacks, various defense strategies have been proposed. A common approach is adversarial training (Madry et al., 2017), where models are trained on adversarial examples generated from a specific attack method. In addition, to improve robustness, most existing defenses require modifications to the original model architecture (Panousis et al., 2021; Liu et al., 2024; Mohammed et al., 2024; Zhou et al., 2023), introducing additional complexity and often being architecture-dependent. Some approaches also involve test-time data purification (Meng & Chen, 2017; Song et al., 2017; Cohen & Giryes, 2024; Tang & Zhang, 2024), which increases latency, limiting their practical applicability.
In this paper, we propose a novel method named Multi-Objective REpresentation Learning (MOREL) that addresses these challenges by focusing on robust feature representation learning. MOREL encourages the model to produce consistent features for inputs within the same class, despite adversarial perturbations. By enhancing the robustness of feature representations, MOREL strengthens the model’s inherent ability to differentiate between classes, making it more resilient to adversarial attacks. The core of our approach is a multi-objective optimization framework that simultaneously optimizes two key objectives: enhancing adversarial robustness and maintaining high classification accuracy. We achieve this by embedding natural and adversarial features into a lower-dimensional space, where cosine similarity and contrastive loss functions are applied to align and tightly cluster these features. The classifier is concurrently motivated to achieve accurate predictions. This approach ensures that the model’s learned representations are robust to adversarial perturbations while preserving the information necessary for accurate classification. The embedding space used during training is discarded, allowing the model to retain its original structure and computational efficiency during inference. This characteristic distinguishes MOREL from many existing defense strategies that either involve architectural changes or rely on additional modules during inference. Through extensive experiments (Sec. 4), we demonstrate that our approach significantly enhances the robustness of DNN models against white-box and black-box adversarial attacks, outperforming existing adversarial training methods that similarly require no architectural changes or test-time data purification, in terms of the accuracy-robustness trade-off. In summary, our key contributions are:
@@ 2.1 Adversarial Training
Adversarial training, introduced by Madry et al. (2017), has emerged as one of the most effective defenses against adversarial attacks. The core idea involves augmenting the training data with adversarial examples generated using methods like Projected Gradient Descent (PGD). While standard adversarial training has proven effective against known attacks, it often results in models becoming overly specialized to the specific types of adversarial examples used during training (Tsipras et al., 2018). To address this limitation, several variants of adversarial training have been proposed. Kannan et al. (2018) introduced Adversarial Logit Pairing (ALP), which enhances robustness by pairing logits from adversarial and clean examples during training. Building on this, they proposed Clean Logit Pairing (CLP), which further refines the approach by focusing specifically on randomly selected clean training examples. Ding et al. (2018) advanced the field with Max-Margin Adversarial (MMA) training, which pushes decision boundaries further from data points, thereby offering enhanced robustness. The TRADES method by Zhang et al. (2019) marked a significant leap forward by explicitly balancing the trade-off between robustness and accuracy through a regularized loss function that minimizes the Kullback-Leibler divergence between predictions on natural and adversarial examples. This was further refined by MART Wang et al. (2019), which focuses on the robustness of misclassified examples, addressing vulnerabilities near the decision boundary. Despite these advancements, common limitations persist, including the challenge of maintaining strong robustness while achieving high accuracy on clean data. Building on these state-of-the-art adversarial training methods, our approach, MOREL, addresses these challenges by strengthening the robustness of DNNs through a robust feature representation learning technique. By considering a multi-objective optimization framework, MOREL aims to achieve the best possible trade-offs between robustness and accuracy—an aspect that, to our knowledge, has not been fully explored in previous work.
@@ 2.2 Insights from Contrastive Learning
To enhance the learning of robust features in the context of adversarial training, our method also draws insights from recent advances in contrastive learning. Contrastive learning has been shown to be effective in producing robust and well-structured feature representations by encouraging similar samples to be closer in the embedding space while pushing dissimilar samples apart (Chen et al., 2020; Gidaris et al., 2018; He et al., 2020). Specifically, Khosla et al. (2020) extend the principles of contrastive learning to a supervised setting. This method leverages label information to group similar examples (i.e., those sharing the same class label) closer together in the feature space. This work informs the design of our embedding space in MOREL, where we apply a multi-positive contrastive loss function (Khosla et al., 2020; Tian et al., 2024) to align natural and adversarial features. By doing so, MOREL not only enhances robustness against adversarial attacks but also ensures that the learned features are tightly clustered and well-separated across different classes, improving both robustness and accuracy.
@@ 3 Methods
We consider a supervised classification problem where a DNN model $f$ parameterized by $\theta\in\Omega$ learns to map an input image $x\in\mathbb{R}^{d}$ to a target class $f(x)=y\in\{1,...,c\}$ where $c\in\mathbb{N}$. An adversarial example $x^{\prime}\in\mathbb{R}^{d}$ is an image obtained by adding a small perturbation of at most $\epsilon>0$ to every pixel of $x$ such that $f(x)\neq f(x^{\prime})$. With a given $l_{p}$-based adversarial region $\mathcal{R}_{p}(x,\epsilon)=\{x^{\prime}\in\mathbb{R}^{d}\ \mid\ \left\|x^{\prime}-x\right\|_{p}\leq\epsilon\}$, the aim of adversarial training (Madry et al., 2017) is typically to approximately minimize the risk on the data distribution $\mathcal{D}$ over adversarial examples:
$\min_{\theta}\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\max_{x^{\prime}\in\mathcal{R}_{p}(x,\epsilon)}\mathcal{L}(\theta,f(x^{\prime}),y)\right]$ (1)
where $\mathcal{L}$ is the loss function. The approximate solutions to the inner maximization problem are derived using a specific attack method to generate adversarial examples, while the outer minimization problem involves training on these generated examples.
3.1 Adversarial Attacks
To generate adversarial examples for training, we use the Projected Gradient Descent (PGD) attack (Madry et al., 2017). It is an iterative method that generates adversarial examples by iteratively applying small perturbations to the input. Given an input image $x$, the true label $y$, a loss function $\mathcal{L}(\theta,x,y)$, and a model parameterized by $\theta$, the PGD attack generates an adversarial example $x^{\prime}$ through the following iterative process for a predefined number of iterations:
@@ 3.1 Adversarial Attacks
where, ${x^{\prime}}^{i}$ is the adversarial example at the $i$-th iteration, $\eta$ the step size, $\epsilon$ the maximum perturbation allowed, and $\text{Proj}_{\mathcal{R}(x,\epsilon)}$ the projection operator that ensures the adversarial example remains within the $\epsilon$-ball centered at $x$. Especially, we consider the $l_{\infty}$-based adversarial region:
$\mathcal{R}(x,\epsilon)=\{x^{\prime}\in\mathbb{R}^{d}\ \mid\ \left\|x^{\prime}-x\right\|_{\infty}\leq\epsilon\}.$
@@ 3.2 Multi-Objective Representation Learning
Training a robust model often results in a decrease in test accuracy. The goal of adversarial robustness is then to mitigate the trade-off between accuracy and robustness, thereby enhancing the model’s performance on both natural and adversarial examples (Zhang et al., 2019; Raghunathan et al., 2020). We approach this challenge as a multi-objective optimization problem. The first objective is to constrain the model to produce features that are as similar as possible for input images within the same class, and as dissimilar as possible from feature distributions of other classes, despite perturbations. The second objective is to enhance the model’s accuracy. We denote the model encoder as $g$ (typically the model without its final layer) and the classifier as $h$ (typically the final layer). Let =$\mathcal{B}=\big{\{}x_{i}\in\mathbb{R}^{d}\mid i\in\{1,...,n\}\big{\}}$ be a batch of $n$ natural images with classes $\big{\{}y_{i}\in\{1,...,c\}\mid i\in\{1,...,n\}\big{\}}=\mathcal{Y}$, and =$\mathcal{B}^{\prime}=\big{\{}x_{i}^{\prime}\in\mathcal{R}_{p}(x_{i},\epsilon)\mid i\in\{1,...,n\}\big{\}}$ its adversarial batch. The encoder then produces features [^1] We use the matrix notation $(z_{i})_{i=1}^{n}=Z\in\mathbb{R}^{n\times o}$, where $Z$ is the concatenation of the $n$ vectors $z_{i}$, each of dimension $o$. In the remainder of this paper, when we use the $\bigoplus$ operation over these matrices, it refers to a concatenation operation.:
$g(\mathcal{B})=(z_{i})_{i=1}^{n},\ \text{and}\ g(\mathcal{B}^{\prime})=(z^{\prime}_{i})_{i=1}^{n}.$ (4)
@@ 3.2.1 Embedding Space with Class-Adaptive Multi-Head Attention
During training, we consider an embedding space that includes a linear layer $L_{\text{e}}$ to project the features from the encoder into a lower-dimensional space:
$L_{\text{e}}((z_{i})_{i=1}^{n})=(s_{i})_{i=1}^{n},\ \text{and}\ L_{\text{e}}((z^{\prime}_{i})_{i=1}^{n})=(s^{\prime}_{i})_{i=1}^{n}.$ (5)
@@
$(s_{i})_{i=1}^{n}=\bigoplus\limits_{y\in\{1,...,c\}}(s_{i}^{y})_{i=1}^{n_{y}},\ \text{and}\ (s^{\prime}_{i})_{i=1}^{n}=\bigoplus\limits_{y\in\{1,...,c\}}({s^{\prime}}_{i}^{y})_{i=1}^{n_{y}}.$ (6)
where $n_{y}$ is the number of features of class $y$ present.
@@
Additionally, a class-adaptive multi-head attention module $M_{\text{e}}$ enables interaction within each lower-dimensional feature group, resulting in richer feature representations. This module functions similarly to the multi-head attention mechanism in the vision transformer (Dosovitskiy et al., 2020; Xiong et al., 2020), where the linearly embedded image patches can be viewed as a lower-dimensional feature group. The key distinction is that our multi-head attention module operates on features from different images (instead of features from the patches of the same image), and we omit any positional embedding mechanism since the position of a feature within its lower-dimensional feature group is irrelevant in our case (otherwise, this would imply keeping track of the position of an image within its batch).
More precisely, given a lower-dimensional feature group $(s_{i}^{y})_{i=1}^{n_{y}}\ \left(\text{or}\ ({s^{\prime}}_{i}^{y})_{i=1}^{n_{y}}\right)$, the module $M_{\text{e}}$ produces the final embedded feature group $(t_{i}^{y})_{i=1}^{n_{y}}\ \left(\text{or}\ ({t^{\prime}}_{i}^{y})_{i=1}^{n_{y}}\right)$ via Algorithm 1.
Algorithm 1 Class-Adaptive Multi-Head Attention
0:  A feature group $(s_{i}^{y})_{i=1}^{n_{y}}=S_{y}\in\mathbb{R}^{n_{y}\times b}$, from class $y$.
0:  The availability of learnable triplet weight matrices $W_{j}^{Q}$, $W_{j}^{K}$ and $W_{j}^{V}\in\mathbb{R}^{b\times b_{j}}\ (b_{j}=b/m)$ for each head $j\in\{1,...,m\}$ as well as an additional learnable weight matrix $W^{O}\in\mathbb{R}^{mb_{j}\times b}$.
1:  for $j=1$ to $m$ do
2:     Normalize $S_{y}$ via layer normalization.
3:     Project $S_{y}$ through linear transformations:$Q_{j}=S_{y}W_{j}^{Q},\ K_{j}=S_{y}W_{j}^{K},\ \text{and}\ V_{j}=S_{y}W_{j}^{V}.$
4:     Get the attention score: $A_{j}={\tt softmax}\left(\dfrac{Q_{j}K_{j}^{T}}{\sqrt{b_{j}}}\right)$
5:     Compute the $j^{\text{th}}$ head output $O_{j}=A_{j}V_{j}$.
6:  end for
7:  Concatenate the outputs from the $m$ heads and project the result through a linear transformation:$O={\tt concat}(O_{1},\dots,O_{m})W^{O}$
8:  return  $(t_{i}^{y})_{i=1}^{n_{y}}=S_{y}+O$
@@
This approach takes advantage of the global context understanding property of the attention mechanism (Dosovitskiy et al., 2020; Han et al., 2022) to capture dependencies and relationships across features within the same group (class). During model evaluation on the test set, the embedding space is discarded, keeping the original model architecture unchanged.
Figure 1: Overview of our proposed MOREL method. During training, encoder $g(\cdot)$ features are projected into a lower-dimensional space using a linear layer $L_{\text{e}}$. A class-adaptive multi-head attention module $M_{\text{e}}$ then facilitates interaction within each of the obtained feature groups. The natural and adversarial embedded features are aligned using $\mathcal{L}_{cosine}$ (8), and $\mathcal{L}_{csl}$ (9) ensures tight clustering of the class-informed embedded features. Simultaneously, the classifier $h(\cdot)$ is encouraged to make accurate predictions. During evaluation, the embedding space is discarded, preserving the original model architecture.
3.2.2 Multi-Objective Optimization

### v4

# Enhancing Adversarial Robustness through Multi-Objective Representation Learning

@@ Abstract
Deep neural networks (DNNs) are vulnerable to small adversarial perturbations, which are tiny changes to the input data that appear insignificant but cause the model to produce drastically different outputs. Many defense methods require modifying model architectures during evaluation or performing test-time data purification. This not only introduces additional complexity but is often architecture-dependent. We show, however, that robust feature learning during training can significantly enhance DNN robustness. We propose MOREL, a multi-objective approach that aligns natural and adversarial features using cosine similarity and multi-positive contrastive losses to encourage similar features for same-class inputs. Extensive experiments demonstrate that MOREL significantly improves robustness against both white-box and black-box attacks. Our code is available at https://github.com/salomonhotegni/MOREL.
Keywords:
Adversarial robustness Multi-objective optimization.
@@ 1 Introduction
Deep neural networks (DNNs) have achieved impressive results in many vision tasks, but their ability to generalize beyond the training distribution remains a major challenge, especially when deployed in safety-critical domains such as autonomous driving and medical diagnosis [2, 20]. In particular, extensive research has demonstrated that DNNs can be fooled by adversarial examples: inputs that differ only by imperceptible perturbations yet cause high-confidence misclassifications [21]. To mitigate the risks posed by adversarial attacks, various defense strategies have been proposed. A common approach is adversarial training [18], where models are trained on adversarial examples generated from a specific attack method. In addition, to improve robustness, most existing defenses require modifications to the original model architecture during evaluation [23, 17], introducing additional complexity and often being architecture-dependent. Some approaches also involve test-time data purification [19, 26], which increases latency, limiting their practical applicability. In this paper, we propose a novel method named Multi-Objective REpresentation Learning (MOREL) that addresses these challenges by focusing on robust feature representation learning. MOREL encourages the model to produce consistent features for inputs within the same class, despite adversarial perturbations. By enhancing the robustness of feature representations, MOREL strengthens the model’s inherent ability to differentiate between classes, making it more resilient to adversarial attacks. The core of our approach is a multi-objective optimization framework that simultaneously optimizes two key objectives: enhancing adversarial robustness and maintaining high classification accuracy. Through extensive experiments (Sec. 4), we demonstrate that our approach significantly enhances the robustness of DNN models against white-box and black-box adversarial attacks in terms of the accuracy-robustness trade-off, outperforming existing adversarial training methods that similarly require no architectural changes during evaluation or test-time data purification. In summary, our key contributions are:
@@ 2.1 Adversarial Training
Adversarial training, introduced by [18], has emerged as one of the most effective defenses against adversarial attacks. The core idea involves augmenting the training data with adversarial examples generated using methods like Projected Gradient Descent (PGD). While standard adversarial training has proven effective against known attacks, it often results in models becoming overly specialized to the specific types of adversarial examples used during training [28]. To address this limitation, several variants of adversarial training have been proposed. [12] introduced Adversarial Logit Pairing (ALP), which enhances robustness by pairing logits from adversarial and clean examples during training. Building on this, they proposed Clean Logit Pairing (CLP), which further refines the approach by focusing specifically on randomly selected clean training examples. The TRADES method by [32] marked a significant leap forward by explicitly balancing the trade-off between robustness and accuracy through a regularized loss function that minimizes the Kullback-Leibler divergence between predictions on natural and adversarial examples. This was further refined by MART [29], which focuses on the robustness of misclassified examples, addressing vulnerabilities near the decision boundary. Despite these advancements, common limitations persist, including the challenge of maintaining strong robustness while achieving high accuracy on clean data. Building on these state-of-the-art adversarial training methods, our approach, MOREL, addresses these challenges by strengthening the robustness of DNNs through a robust feature representation learning technique. By considering a multi-objective optimization framework, MOREL aims to achieve the best possible trade-offs between robustness and accuracy, an aspect that, to our knowledge, has not been fully explored in previous work.
@@ 2.2 Insights from Contrastive Learning
To enhance the learning of robust features in the context of adversarial training, our method also draws insights from recent advances in contrastive learning. Contrastive learning has been shown to be effective in producing robust and well-structured feature representations by encouraging similar samples to be closer in the embedding space while pushing dissimilar samples apart [4]. Specifically, [14] extend the principles of contrastive learning to a supervised setting. This method leverages label information to group similar examples (i.e., those sharing the same class label) closer together in the feature space. This work informs the design of our embedding space in MOREL, where we apply a multi-positive contrastive loss function [14, 27] to align natural and adversarial features. By doing so, MOREL not only enhances robustness against adversarial attacks but also ensures that the learned features are tightly clustered and well-separated across different classes, improving both robustness and accuracy.
@@ 3 Methods
We consider supervised classification problems where a DNN model $f$ parameterized by $\theta\in\Omega$ learns to map an input image $x\in\mathbb{R}^{d}$ to a target class $f(x)=y\in\{1,...,c\}$ where $c\in\mathbb{N}$. An adversarial example $x^{\prime}\in\mathbb{R}^{d}$ is an image obtained by adding imperceptible perturbations to $x$ such that $f(x)\neq f(x^{\prime})$. With a given $l_{p}$-based adversarial region $\mathcal{R}_{p}(x,\epsilon)=\{x^{\prime}\in\mathbb{R}^{d}\ \mid\ \left\|x^{\prime}-x\right\|_{p}\leq\epsilon\}$ and a loss function $\mathcal{L}$, the aim of adversarial training [18] is typically to approximately minimize the risk on the data distribution $\mathcal{D}$ over adversarial examples: $\min_{\theta}\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\max_{x^{\prime}\in\mathcal{R}_{p}(x,\epsilon)}\mathcal{L}(\theta,f(x^{\prime}),y)\right].$ The approximate solutions to the inner maximization problem are derived using a specific attack method to generate adversarial examples, while the outer minimization problem involves training on these generated examples. To generate adversarial examples for training, we use the Projected Gradient Descent (PGD) attack [18]. PGD is a method that generates adversarial examples by iteratively applying small perturbations to the input. Given an input image $x$, the true label $y$, a loss function $\mathcal{L}(\theta,x,y)$, and a model parameterized by $\theta$, the PGD attack generates an adversarial example $x^{\prime}$ through the following iterative process for a predefined number of iterations:
@@
where, ${x^{\prime}}^{i}$ is the adversarial example at the $i$-th iteration, $\eta$ the step size, $\epsilon$ the maximum perturbation allowed, and $\text{Proj}_{\mathcal{R}(x,\epsilon)}$ the projection operator that ensures the adversarial example remains within the $\epsilon$-ball centered at $x$. We consider the $l_{\infty}$-based adversarial region: $\mathcal{R}(x,\epsilon)=\{x^{\prime}\in\mathbb{R}^{d}\ \mid\ \left\|x^{\prime}-x\right\|_{\infty}\leq\epsilon\}.$
@@ 3.1 Multi-Objective Representation Learning
Training a robust model often results in a decrease in test accuracy. The goal of adversarial robustness is then to mitigate the trade-off between accuracy and robustness, thereby enhancing the model’s performance on both natural and adversarial examples [32, 25]. We approach this challenge as a multi-objective optimization problem. The first objective is to constrain the model to produce features that are as similar as possible for input images within the same class, and as dissimilar as possible from feature distributions of other classes, despite perturbations. The second objective is to enhance the model’s accuracy. We denote the model encoder as $g$ (typically the model without its final layer) and the classifier as $h$ (typically the final layer). Let =$\mathcal{B}=\big{\{}x_{i}\in\mathbb{R}^{d}\mid i\in\{1,...,n\}\big{\}}$ be a batch of $n$ natural images with classes $\big{\{}y_{i}\in\{1,...,c\}\mid i\in\{1,...,n\}\big{\}}=\mathcal{Y}$, and =$\mathcal{B}^{\prime}=\big{\{}x_{i}^{\prime}\in\mathcal{R}_{p}(x_{i},\epsilon)\mid x_{i}\in\mathcal{B}\big{\}}$ its adversarial batch. The encoder then produces features [^1] We use the matrix notation $(z_{i})_{i=1}^{n}=Z\in\mathbb{R}^{n\times o}$, where $Z$ is the concatenation of the $n$ vectors $z_{i}$, each of dimension $o$. :
$g(\mathcal{B})=(z_{i})_{i=1}^{n}\in\mathbb{R}^{n\times o},\ \text{and}\ g(\mathcal{B}^{\prime})=(z^{\prime}_{i})_{i=1}^{n}\in\mathbb{R}^{n\times o}.\vspace{-10pt}$ (3)
@@ 3.1.1 Embedding Space with Class-Adaptive Multi-Head Attention
During training, we consider an embedding space that includes a linear layer $L_{\text{e}}$ of size $b$ to project the features from the encoder into a lower-dimensional space:
$L_{\text{e}}((z_{i})_{i=1}^{n})=(s_{i})_{i=1}^{n}\in\mathbb{R}^{n\times b},\ \text{and}\ L_{\text{e}}((z^{\prime}_{i})_{i=1}^{n})=(s^{\prime}_{i})_{i=1}^{n}\in\mathbb{R}^{n\times b}.\vspace{-5pt}$ (4)
@@
$(s_{i})_{i=1}^{n}=\bigoplus\limits_{y\in\{1,...,c\}}(s_{i}^{y})_{i=1}^{n_{y}},\ \text{and}\ (s^{\prime}_{i})_{i=1}^{n}=\bigoplus\limits_{y\in\{1,...,c\}}({s^{\prime}}_{i}^{y})_{i=1}^{n_{y}}.\vspace{-5pt}$ (5)
@@
where $n_{y}$ is the number of features of class $y$ present within the batch, and $\bigoplus$ refers to a concatenation operation. Additionally, a class-adaptive multi-head attention module $M_{\text{e}}$ enables interaction within each lower-dimensional feature group, resulting in richer feature representations. This module functions similarly to the multi-head attention mechanism in the vision transformer (ViT) [6], where the linearly embedded image patches can be viewed as a lower-dimensional feature group. The key distinction is that our multi-head attention module operates on features from different images (instead of features from the patches of the same image), and we omit any positional embedding mechanism since the position of a feature within its lower-dimensional feature group is irrelevant in our case (otherwise, this would imply keeping track of the position of an image within its batch). More precisely, given a lower-dimensional feature group $(s_{i}^{y})_{i=1}^{n_{y}}\in\mathbb{R}^{n_{y}\times b}\ \left(\text{or}\ ({s^{\prime}}_{i}^{y})_{i=1}^{n_{y}}\right)$, the module $M_{\text{e}}$ produces the final embedded feature group $(t_{i}^{y})_{i=1}^{n_{y}}\in\mathbb{R}^{n_{y}\times b}\ \left(\text{or}\ ({t^{\prime}}_{i}^{y})_{i=1}^{n_{y}}\right)$ via Algorithm 1. This results in a total cost of $\sum_{y}O\bigl{(}n_{y}^{2}\times b\bigr{)}$, which in the worst case is $O(n^{2}\times b)$ per forward pass. With a moderate batch size $n$, this quadratic cost remains manageable. All such groups are concatenated back along the batch dimension to form:
$T=\bigoplus\limits_{y\in\{1,...,c\}}(t_{i}^{y})_{i=1}^{n_{y}}\in\mathbb{R}^{n\times b},\ \text{and}\ T^{\prime}=\bigoplus\limits_{y\in\{1,...,c\}}({t^{\prime}}_{i}^{y})_{i=1}^{n_{y}}\in\mathbb{R}^{n\times b}.$ (6)
@@
This approach takes advantage of the global context understanding property of the attention mechanism [6, 9] to capture dependencies and relationships across features within the same group (class). During model evaluation on the test set, the embedding space is discarded, keeping the original model architecture unchanged.
Algorithm 1 Class-Adaptive Multi-Head Attention 0:  A feature group $(s_{i}^{y})_{i=1}^{n_{y}}=S_{y}\in\mathbb{R}^{n_{y}\times b}$, from class $y$. 0:  The availability of learnable triplet weight matrices $W_{j}^{Q}$, $W_{j}^{K}$ and $W_{j}^{V}\in\mathbb{R}^{b\times b_{j}}\ (b_{j}=b/m)$ for each head $j\in\{1,...,m\}$ as well as an additional learnable weight matrix $W^{O}\in\mathbb{R}^{mb_{j}\times b}$. 1:  for $j=1$ to $m$ do 2:     Normalize $S_{y}$ via layer normalization. 3:     Project $S_{y}$ through linear transformations:$Q_{j}=S_{y}W_{j}^{Q},\ K_{j}=S_{y}W_{j}^{K},\ \text{and}\ V_{j}=S_{y}W_{j}^{V}.$ 4:     Get the attention score: $A_{j}={\tt softmax}\left(\dfrac{Q_{j}K_{j}^{T}}{\sqrt{b_{j}}}\right)$ 5:     Compute the $j^{\text{th}}$ head output $O_{j}=A_{j}V_{j}$. 6:  end for 7:  Concatenate the outputs from the $m$ heads and project the result through a linear transformation: $O={\tt concat}(O_{1},\dots,O_{m})W^{O}$ 8:  Output: $(t_{i}^{y})_{i=1}^{n_{y}}=S_{y}+O$
3.1.2 Multi-Objective Optimization (MOO)